# Enumerating log rational curves on some toric varieties

### v2

@@ 1.1. Counting curves and log curves
Let $X$ be a smooth, projective variety. Gromov-Witten invariants of $X$ are intersection numbers against the virtual fundamental class [¯$[\overline{\mathcal{M}}_{g,n}(X,\beta)]^{\text{vir}}$ on the moduli space of stable maps, and virtually enumerate curves $C$ on $X$ [Beh97, BF97].
Degeneration of the target variety has long played an important role in Gromov-Witten theory [Li01, Li02]. In the simplest situation, $X$ degenerates into the union of two smooth, projective varieties $X_{1}\cup X_{2}$ meeting along a smooth divisor $D$. One hopes to count curves on $X$ in terms of counts of curves of $X_{1},X_{2}$, but must now keep track of the incidences of the curves on $X_{j}$ with the divisor $D$. This leads naturally to the theory of logarithmic Gromov-Witten invariants of the log scheme $(X_{j},D)$ [GS13, AC14a, AC14b].
Calculations of Gromov-Witten invariants where the (pointed) complex structure of the domain curve is fixed also drew much attention early in the subject. The Vafa-Intriligator formula [BDW96, ST97, MO07] determines Gromov-Witten invariants of Grassmannians upon restriction over a fixed point of ¯$\overline{\mathcal{M}}_{g,n}$ [MOP11]. After work of Tevelev [Tev24], the fixed-domain Gromov-Witten invariants have received renewed interest. The “virtual Tevelev degrees” of $X$ are defined by Buch-Pandharipande [BP21] to be the virtual degrees of the forgetful morphisms
@@
A feature of the fixed-domain setting which has motivated much of the more recent work is that, compared to arbitrary Gromov-Witten invariants, the fixed-domain invariants tend to be enumerative much more often [LP25, BLL+24, CD25]. Roughly speaking, this means that the virtual invariants compute the actual degree of $\tau$, denoted $\mathrm{Tev}^{X}_{g,n,\beta}$, upon restricting to the open locus of maps $\mathcal{M}_{g,n}(X,\beta)$ with smooth domain. Concretely, $\mathrm{Tev}^{X}_{g,n,\beta}$ is the number of maps $f:C\to X$ in class $\beta$ from a fixed, general curve $C$ to $X$ satisfying $n$ incidence conditions $f(p_{i})=x_{i}$ at general points of $C$ and $X$. Moreover, even when the enumerativity fails, the geometric (non-virtual) counts $\mathrm{Tev}^{X}_{g,n,\beta}$ can sometimes be accessed using different techniques [CPS22, FL23, CL23b, CD24, CL25b, Lia24, CL23a, CL25a].
For toric varieties $X$, Cela–Iribar López [CIL24] have considered the logarithmic, fixed-domain version of the problem, where the tangency profile of $f$ along the toric boundary of $X$ is constrained. For domain curves of genus 0, they have proven a tropical correspondence theorem [CIL24, Theorem 5], showing that the (geometric or virtual) count of log rational curves on $X$ with the maximal number of incidence conditions imposed at general points matches a purely combinatorial enumeration of tropical curves. In fact, a more general tropical correspondence theorem for log Gromov-Witten invariants of toric varieties in any genus and with arbitrary $\psi$-class insertions (descendants) had been obtained earlier by Mandel-Ruddat [MR20]. The fixed-domain setting may be recovered by inserting appropriate $\psi$-classes.
Using the tropical correspondence theorem, the genus 0, logarithmic, fixed-domain invariants for Hirzebruch surfaces are computed [CIL24, Theorem 7], and conjectures for other toric varieties are proposed [CIL24, Conjectures 14-15]. See also [Zho25, GRZZ25] for further recent work exploiting the connection between log Gromov-Witten theory of projective bundles on toric varieties and tropical curve-counting. The purpose of the present paper is to address the conjectures of Cela–Iribar López, but in the realm of algebro-geometric intersection theory, rather than tropical geometry. Our approach via moduli spaces of “naive log quasimaps” is direct, avoiding the combinatorial analyses typical in tropical enumerations, and leads to explicit formulas. The invariants of interest are reviewed in the next section.
@@ 1.2. Logarithmic Tevelev degrees
Thus, the points $q_{\rho,v}$ are exactly those points mapping to the toric boundary of $X$, with multiplicities $\mu_{\rho,v}$. In particular, $f$ maps the $p_{i}$ to $X^{\circ}$, and $f_{*}[C]=\beta$. The letter $\Gamma$ is used to denote the combinatorial data of the curve class $\beta$, the integer $n$, and the vectors $\mu_{\rho}$ specifiying the tangency profile of $f$ along the boundary.
@@
Let $\mathcal{M}_{\Gamma}(X)$ be the moduli space of pointed maps $f$ as above.
$\mathcal{M}_{\Gamma}(X)$ is a smooth, quasi-projective variety of dimension $(m+n-3)+\dim(X)$. The forgetful morphism
$\tau:\mathcal{M}_{\Gamma}(X)\to\mathcal{M}_{0,n}\times X^{n}$
@@
(cf. [CIL24, Definition 3]) Assuming (1), the (genus 0) logarithmic Tevelev degree $\mathrm{logTev}^{X}_{\Gamma}$ is defined by $\deg(\tau)$.
@@
The logarithmic Tevelev degree $\mathrm{Tev}^{X}_{\Gamma}$ in [CIL24] is instead defined by the quotient of $\deg(\tau)$ by the combinatorial factor
@@
The moduli space of genus 0, log stable maps ¯$\overline{\mathcal{M}}_{\Gamma}(X)$ is irreducible of expected dimension [Ran17, Proposition 3.3.5], so one can just as well define $\mathrm{logTev}^{X}_{\Gamma}$ to be the degree of
$\tau:\overline{\mathcal{M}}_{\Gamma}(X)\to\overline{\mathcal{M}}_{0,n}\times X^{n}.$
In this way, $\mathrm{logTev}^{X}_{\Gamma}$ is given by the integral of a tautological class on ¯$\overline{\mathcal{M}}_{\Gamma}(X)$. In higher genus, the moduli space of log stable maps no longer has expected dimension, so one should instead integrate against the virtual fundamental class. We work in this paper only in genus 0, where the virtual class is the usual fundamental class.
@@ 1.3. New results
Theorem 1.3.1 appears as Theorem 4.2.1 in the main body of the paper. The torus-invariant divisors $D_{j}\subset X_{r,s,a}$ are indexed by the integers $j=0,\ldots,r+s+1$, see §2.1. The integer $m_{0}$ appearing in the binomial coefficient of the main formula corresponds in particular to the divisor $D_{0}$, which in turn is the vanishing locus of the coordinate corresponding to the summand $\mathcal{O}(-a)$ in $X_{r,s,a}$. The $s=1$ case of Theorem 1.3.1 proves [CIL24, Conjecture 14]. See §4.3 for other specializations of Theorem 1.3.1.
Our proof of Theorem 1.3.1 proceeds by an intersection-theoretic calculation on a moduli space $Q_{\Gamma}(X_{r,s,a})$ of “naive log quasimaps,” in which $\mathcal{M}_{\Gamma}(X_{r,s,a})$ is a dense open subset, see §2.3. These moduli spaces are similar to those considered in [CL23a, CL25a] for blow-ups of ${\mathbb{P}}^{r}$, but incorporate the data of tangency profiles along the toric boundary.
A subtlety is that, unlike on the moduli space of log stable maps ¯$\overline{\mathcal{M}}_{\Gamma}(X_{r,s,a})$, the natural intersection numbers on $Q_{\Gamma}(X_{r,s,a})$ are not always enumerative. We prove, however, that in all cases where the enumerativity fails, the logarithmic Tevelev degree is zero. The advantage of the space $Q_{\Gamma}(X_{r,s,a})$ is that it admits a simple description as a tower of projective bundles over a product of projective lines, so the intersection-theoretic calculation is relatively straightforward.
@@
The other $X$ we consider is the blow-up of ${\mathbb{P}}^{r}$ at $r$ (torus-fixed) points, which occupies §5. While we do not determine the logarithmic Tevelev degrees completely, we show in Theorem 5.2.2 that the prediction of [CIL24, Conjecture 15] does not always hold, already when $r=2$. We follow a similar program to the case $X=X_{r,s,a}$, constructing a moduli space $Q_{\Gamma}(X)$ of naive log quasimaps to $X$. While the value of $\mathrm{logTev}^{X}_{\Gamma}$ predicted by [CIL24, Conjecture 15] equals a tautological intersection number on our moduli space $Q_{\Gamma}(X)$, we find that this intersection number can fail to be enumerative even when $\mathrm{logTev}^{X}_{\Gamma}$ cannot be proven to be zero. Using the excess intersection formula, we compute $\mathrm{logTev}^{X}_{\Gamma}$ to be both non-zero and different from the value predicted by [CIL24, Conjecture 15] in one example (and the method extends to many more).
@@ 1.5. Acknowledgments
Preliminary discussions related to this project took place at the SLMath (formerly MSRI) summer school on Algebraic Curves in July 2024, and at the workshop “Tevelev Degrees and related topics” at UIUC in October 2024. We thank the organizers and other participants of these events for providing an inspiring work environment. We also thank Alessio Cela and Helge Ruddat for comments on drafts of this work. C.L. has been supported by NSF Postdoctoral Fellowship DMS-2001976, an AMS-Simons travel grant, and the Summer Scholars program at Tufts University.
@@ 2.1. The target variety
Furthermore, the collections of coordinates $\{y_{1},\ldots,y_{r+1}\}$ and $\{y_{0},y_{r+2},\ldots,y_{r+s+1}\}$ are required not all to vanish. The point $\pi(x)\in{\mathbb{P}}^{r}$ is given by $[y_{1}:\cdots:y_{r+1}]$, and the fiber of $\pi$ over $[y_{1}:\cdots:y_{r+1}]$ is given by the projective space with coordinates $y_{0},y_{r+2},\ldots,y_{r+s+1}$.
$X_{r,s,a}$ is toric, with action (×$(\mathbb{C}^{\times})^{r+s}\times X_{r,s,a}\to X_{r,s,a}$ given by
@@ 2.2. Maps to $X_{r,s,a}$
In this section, we discuss the parametrization of maps from rational curves to $X_{r,s,a}$ with prescribed tangency orders at the boundary. See [Cox95] for a description of maps to any toric variety. Fix data $\Gamma$ (see §1.2) associated to $X_{r,s,a}$, and let $f:{\mathbb{P}}^{1}\to X_{r,s,a}$ be a map in $\mathcal{M}_{\Gamma}(X_{r,s,a})$. The map $f$ fits into a commutative diagram
@@
The map $g:{\mathbb{P}}^{1}\to{\mathbb{P}}^{r}$ is given by a tuple $[g_{1}:\cdots:g_{r+1}]$, where $g_{j}\in H^{0}({\mathbb{P}}^{1},\mathcal{O}(b))$, for some $b\geq 0$. If $g$ is fixed, then $f$ is determined by the data of a surjection
@@ 2.3. Naive log quasimaps to $X_{r,s,a}$
and let $\nu:{\mathbb{P}}^{1}\times\mathcal{B}\longrightarrow Y$ be the projection. $\mathcal{B}$ is simply a product of copies of ${\mathbb{P}}^{1}$, parametrizing possible positions of the variable points $q_{j,v}$.
For each $j\in[0,r+s+1]$ and $v\in[1,m_{j}]$, let $\mathcal{D}_{j,v}\subset{\mathbb{P}}^{1}\times Y$ be the pullback of the diagonal from ${\mathbb{P}}^{1}\times{\mathbb{P}}_{j,v}$. For each $j$, define the Cartier divisor
@@
We refer to $Q_{\Gamma}(X_{r,s,a})$ as the moduli space of naive log quasimaps to $X_{r,s,a}$. Ignoring the marked points $q_{j,v}$, elements $f\in Q_{\Gamma}(X_{r,s,a})$ are indeed quasimaps in the sense of Ciocan-Fontanine–Kim [CFK10]. If $f\in Q_{\Gamma}(X_{r,s,a})$ is not bpf, then it still defines a rational map ${\mathbb{P}}^{1}\dashrightarrow X_{r,s,a}$. After twisting down base-points, this map extends canonically to an actual map, but no longer in class $\beta$. A version of this twisting is made explicit in §3.2.
The adjective “log” refers to the fact that the data of the sections $g_{j}$ underlying $f$ are considered relative to the data of underlying divisors, even when the $g_{j}$ are identically zero. Our space is “naive” in the sense that we allow the points $q_{j,v}$ to collide arbitrarily, rather than requring the data of a stable, logarithmic curve. The actual moduli space of log quasimaps to more general targets is constructed in [Sha24].

### v3

@@ 1.1. Counting curves and log curves
Let $X$ be a smooth, projective variety. Gromov-Witten invariants of $X$ are intersection numbers against the virtual fundamental class $\overline{\mathcal{M}}_{g,n}(X,\beta)]^{\text{vir}}$ on the moduli space of stable maps, and virtually enumerate curves $C$ on $X$ [behrend, bf].
Degeneration of the target variety has long played an important role in Gromov-Witten theory [li1, li2]. In the simplest situation, $X$ degenerates into the union of two smooth, projective varieties $X_{1}\cup X_{2}$ meeting along a smooth divisor $D$. One hopes to count curves on $X$ in terms of counts of curves of $X_{1},X_{2}$, but must now keep track of the incidences of the curves on $X_{j}$ with the divisor $D$. This leads naturally to the theory of logarithmic Gromov-Witten invariants of the log scheme $(X_{j},D)$ [gs, ac1, ac2].
Calculations of Gromov-Witten invariants where the (pointed) complex structure of the domain curve is fixed also drew much attention early in the subject. The Vafa-Intriligator formula [bdw, st, mo] determines Gromov-Witten invariants of Grassmannians upon restriction over a fixed point of $\overline{\mathcal{M}}_{g,n}$ [mop]. After work of Tevelev [tevelev], the fixed-domain Gromov-Witten invariants have received renewed interest. The “virtual Tevelev degrees” of $X$ are defined by Buch-Pandharipande [bp] to be the virtual degrees of the forgetful morphisms
@@
A feature of the fixed-domain setting which has motivated much of the more recent work is that, compared to arbitrary Gromov-Witten invariants, the fixed-domain invariants tend to be enumerative much more often [lp, bllrst, celadoan]. Roughly speaking, this means that the virtual invariants compute the actual degree of $\tau$, denoted $\mathrm{Tev}^{X}_{g,n,\beta}$, upon restricting to the open locus of maps $\mathcal{M}_{g,n}(X,\beta)$ with smooth domain. Concretely, $\mathrm{Tev}^{X}_{g,n,\beta}$ is the number of maps $f:C\to X$ in class $\beta$ from a fixed, general curve $C$ to $X$ satisfying $n$ incidence conditions $f(p_{i})=x_{i}$ at general points of $C$ and $X$. Moreover, even when the enumerativity fails, the geometric (non-virtual) counts $\mathrm{Tev}^{X}_{g,n,\beta}$ can sometimes be accessed using different techniques [cps, fl, cl1, cavdaw, cl_hirzebruch, lian_pr, cl2, cl_complete].
For toric varieties $X$, Cela–Iribar López [cil] have considered the logarithmic, fixed-domain version of the problem, where the tangency profile of $f$ along the toric boundary of $X$ is constrained. For domain curves of genus 0, they have proven a tropical correspondence theorem [cil, Theorem 5], showing that the (geometric or virtual) count of log rational curves on $X$ with the maximal number of incidence conditions imposed at general points matches a purely combinatorial enumeration of tropical curves. In fact, a more general tropical correspondence theorem for log Gromov-Witten invariants of toric varieties in any genus and with arbitrary $\psi$-class insertions (descendants) had been obtained earlier by Mandel-Ruddat [mr]. The fixed-domain setting may be recovered by inserting appropriate $\psi$-classes.
Using the tropical correspondence theorem, the genus 0, logarithmic, fixed-domain invariants for Hirzebruch surfaces are computed [cil, Theorem 7], and conjectures for other toric varieties are proposed [cil, Conjectures 14-15]. See also [zhou, grzz] for further recent work exploiting the connection between log Gromov-Witten theory of projective bundles on toric varieties and tropical curve-counting. The purpose of the present paper is to address the conjectures of Cela–Iribar López, but in the realm of algebro-geometric intersection theory, rather than tropical geometry. Our approach via moduli spaces of “naive log quasimaps” is direct, avoiding the combinatorial analyses typical in tropical enumerations, and leads to explicit formulas. The invariants of interest are reviewed in the next section.
@@ 1.2. Logarithmic Tevelev degrees
Thus, the points $q_{\rho,v}$ are exactly those points mapping to the toric boundary of $X$, with multiplicities $\mu_{\rho,v}$. In particular, $f$ maps the $p_{i}$ to $X^{\circ}$, and $f_{*}[C]=\beta$. The letter $\Gamma$ is used to denote the combinatorial data of the curve class $\beta$ and the vectors $\mu_{\rho}$ specifiying the tangency profile of $f$ along the boundary. The additional variable $n$ denotes the number of marked points $p_{i}$.
@@
Let $\mathcal{M}_{\Gamma,n}(X)$ be the moduli space of pointed maps $f$ as above. Isomorphisms of maps $f\cong f^{\prime}$ are commutative diagrams
${({\mathbb{P}}^{1},\{p_{i}\}_{i=1}^{n},\{\{q_{\rho,v}\}_{\rho\in\Sigma(1)}\}_{v=1}^{m_{\rho}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${X}$${({\mathbb{P}}^{1},\{p^{\prime}_{i}\}_{i=1}^{n},\{\{q^{\prime}_{\rho,v}\}_{\rho\in\Sigma(1)}\}_{v=1}^{m_{\rho}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f^{\prime}}$
where the vertical arrow is an isomorphism of pointed curves.
$\mathcal{M}_{\Gamma,n}(X)$ is a smooth, quasi-projective variety of dimension $(m+n-3)+\dim(X)$. The forgetful morphism
$\tau:\mathcal{M}_{\Gamma,n}(X)\to\mathcal{M}_{0,n}\times X^{n}$
@@
(cf. [cil, Definition 3]) Assuming (1), the (genus 0) logarithmic Tevelev degree $\mathrm{logTev}^{X}_{\Gamma}$ is defined by $\deg(\tau)$.
@@
The logarithmic Tevelev degree $\mathrm{Tev}^{X}_{\Gamma}$ in [cil] is instead defined by the quotient of $\deg(\tau)$ by the combinatorial factor
@@
The moduli space of genus 0, log stable maps $\overline{\mathcal{M}}_{\Gamma,n}(X)$ is irreducible of expected dimension [ranganathan, Proposition 3.3.5], so one can just as well define $\mathrm{logTev}^{X}_{\Gamma}$ to be the degree of
$\tau:\overline{\mathcal{M}}_{\Gamma,n}(X)\to\overline{\mathcal{M}}_{0,n}\times X^{n}.$
In this way, $\mathrm{logTev}^{X}_{\Gamma}$ is given by the integral of a tautological class on $\overline{\mathcal{M}}_{\Gamma,n}(X)$. In higher genus, the moduli space of log stable maps no longer has expected dimension, so one should instead integrate against the virtual fundamental class. We work in this paper only in genus 0, where the virtual class is the usual fundamental class.
For calculations, it will be useful to consider log maps to ${\mathbb{P}}^{1}$ where the marked points $p_{i}$ are not part of the data. Fix the data of $\beta,\mu_{\rho}$ as above.
Definition 1.2.3.
Let $\mathcal{M}_{\Gamma}(X)$ be the moduli space of pointed maps
$f:({\mathbb{P}}^{1},\{\{q_{\rho,v}\}_{\rho\in\Sigma(1)}\}_{v=1}^{m_{\rho}})\to X$
with $f_{*}[{\mathbb{P}}^{1}]=\beta$ and the same tangency conditions as before. Isomorphisms of maps $f\cong f^{\prime}$ are commutative diagrams
${({\mathbb{P}}^{1},\{\{q_{\rho,v}\}_{\rho\in\Sigma(1)}\}_{v=1}^{m_{\rho}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$$\scriptstyle{\mathrm{id}}$${X}$${({\mathbb{P}}^{1},\{\{q_{\rho,v}\}_{\rho\in\Sigma(1)}\}_{v=1}^{m_{\rho}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f^{\prime}}$
where the vertical arrow is the identity.
That is, $\mathcal{M}_{\Gamma}(X)$ parametrizes log maps with the same tangency profiles as $\mathcal{M}_{\Gamma,n}(X)$, except that the domain curve is viewed as a parametrized ${\mathbb{P}}^{1}$ (as opposed to a genus 0 curve up to isomorphism), and the points $p_{i}$ are not part of the data. $\mathcal{M}_{\Gamma}(X)$ is a smooth, quasi-projective variety of dimension $m+\dim(X)$. We will still be interested in imposing incidence conditions at fixed, general points $p_{i}\in{\mathbb{P}}^{1}$. After fixing such points, in contrast to $\mathcal{M}_{\Gamma,n}(X)$, maps $f\in\mathcal{M}_{\Gamma}(X)$ may send the $p_{i}$ to the boundary of $X$; equivalently, the points $q_{\rho,v}$ may become equal to the $p_{i}$.
@@ 1.3. New results
Theorem 1.3.1 appears as Theorem 4.2.1 in the main body of the paper. The torus-invariant divisors $D_{j}\subset X_{r,s,a}$ are indexed by the integers $j=0,\ldots,r+s+1$, see §2.1. The integer $m_{0}$ appearing in the binomial coefficient of the main formula corresponds in particular to the divisor $D_{0}$, which in turn is the vanishing locus of the coordinate corresponding to the summand $\mathcal{O}(-a)$ in $X_{r,s,a}$. The $s=1$ case of Theorem 1.3.1 proves [cil, Conjecture 14]. See §4.3 for other specializations of Theorem 1.3.1.
Our proof of Theorem 1.3.1 proceeds by an intersection-theoretic calculation on a moduli space $Q_{\Gamma}(X_{r,s,a})$ of “naive log quasimaps,” in which $\mathcal{M}_{\Gamma}(X_{r,s,a})$ (Definition 1.2.3) is a dense open subset, see §2.3. These moduli spaces are similar to those considered in [cl2, cl_complete] for blow-ups of ${\mathbb{P}}^{r}$, but incorporate the data of tangency profiles along the toric boundary.
A subtlety is that, unlike on the moduli space of log stable maps $\overline{\mathcal{M}}_{\Gamma,n}(X_{r,s,a})$, the natural intersection numbers on $Q_{\Gamma}(X_{r,s,a})$ are not always enumerative. We prove, however, that in all cases where the enumerativity fails, the logarithmic Tevelev degree is zero. The advantage of the space $Q_{\Gamma}(X_{r,s,a})$ is that it admits a simple description as a tower of projective bundles over a product of projective lines, so the intersection-theoretic calculation is relatively straightforward.
@@
The other $X$ we consider is the blow-up of ${\mathbb{P}}^{r}$ at $r$ (torus-fixed) points, which occupies §5. While we do not determine the logarithmic Tevelev degrees completely, we show in Theorem 5.2.2 that the prediction of [cil, Conjecture 15] does not always hold, already when $r=2$. We follow a similar program to the case $X=X_{r,s,a}$, constructing a moduli space $Q_{\Gamma}(X)$ of naive log quasimaps to $X$. While the value of $\mathrm{logTev}^{X}_{\Gamma}$ predicted by [cil, Conjecture 15] equals a tautological intersection number on our moduli space $Q_{\Gamma}(X)$, we find that this intersection number can fail to be enumerative even when $\mathrm{logTev}^{X}_{\Gamma}$ cannot be proven to be zero. Using the excess intersection formula, we compute $\mathrm{logTev}^{X}_{\Gamma}$ to be both non-zero and different from the value predicted by [cil, Conjecture 15] in one example (and the method extends to many more).
@@ 1.5. Acknowledgments
Preliminary discussions related to this project took place at the SLMath (formerly MSRI) summer school on Algebraic Curves in July 2024, and at the workshop “Tevelev Degrees and related topics” at UIUC in October 2024. We thank the organizers and other participants of these events for providing an inspiring work environment. We also thank Alessio Cela, Helge Ruddat, and the anonymous referee for comments on drafts of this work. C.L. has been supported by NSF Postdoctoral Fellowship DMS-2001976, an AMS-Simons travel grant, and the Summer Scholars program at Tufts University.
@@ 2.1. The target variety
Furthermore, the collections of coordinates $\{y_{1},\ldots,y_{r+1}\}$ and $\{y_{0},y_{r+2},\ldots,y_{r+s+1}\}$ are required not all to vanish. The point $\pi(x)\in{\mathbb{P}}^{r}$ is given by $y_{1}:\cdots:y_{r+1}]$, and the fiber of $\pi$ over $y_{1}:\cdots:y_{r+1}]$ is given by the projective space with coordinates $y_{0},y_{r+2},\ldots,y_{r+s+1}$.
$X_{r,s,a}$ is toric, with action $(\mathbb{C}^{\times})^{r+s}\times X_{r,s,a}\to X_{r,s,a}$ given by
@@ 2.2. Maps to $X_{r,s,a}$
In this section, we discuss the parametrization of maps from rational curves to $X_{r,s,a}$ with prescribed tangency orders at the boundary. See [cox] for a description of maps to any toric variety. Fix data $\Gamma$ (see §1.2) associated to $X_{r,s,a}$, and let $f:{\mathbb{P}}^{1}\to X_{r,s,a}$ be a map in $\mathcal{M}_{\Gamma}(X_{r,s,a})$ (Definition 1.2.3). The map $f$ fits into a commutative diagram
@@
The map $g:{\mathbb{P}}^{1}\to{\mathbb{P}}^{r}$ is given by a tuple $g_{1}:\cdots:g_{r+1}]$, where $g_{j}\in H^{0}({\mathbb{P}}^{1},\mathcal{O}(b))$, for some $b\geq 0$. If $g$ is fixed, then $f$ is determined by the data of a surjection
@@ 2.3. Naive log quasimaps to $X_{r,s,a}$
and let $\nu:{\mathbb{P}}^{1}\times\mathcal{B}\longrightarrow\mathcal{B}$ be the projection. $\mathcal{B}$ is simply a product of copies of ${\mathbb{P}}^{1}$, parametrizing possible positions of the variable points $q_{j,v}$.
For each $j\in[0,r+s+1]$ and $v\in[1,m_{j}]$, let $\mathcal{D}_{j,v}\subset{\mathbb{P}}^{1}\times\mathcal{B}$ be the pullback of the diagonal from ${\mathbb{P}}^{1}\times{\mathbb{P}}_{j,v}$. For each $j$, define the Cartier divisor
@@
We refer to $Q_{\Gamma}(X_{r,s,a})$ as the moduli space of naive log quasimaps to $X_{r,s,a}$. Ignoring the marked points $q_{j,v}$, elements $f\in Q_{\Gamma}(X_{r,s,a})$ are indeed quasimaps in the sense of Ciocan-Fontanine–Kim [cfk]. If $f\in Q_{\Gamma}(X_{r,s,a})$ is not bpf, then it still defines a rational map ${\mathbb{P}}^{1}\dashrightarrow X_{r,s,a}$. After twisting down base-points, this map extends canonically to an actual map, but no longer in class $\beta$. A version of this twisting is made explicit in §3.2.
The adjective “log” refers to the fact that the data of the sections $g_{j}$ underlying $f$ are considered relative to the data of underlying divisors, even when the $g_{j}$ are identically zero. Our space is “naive” in the sense that we allow the points $q_{j,v}$ to collide arbitrarily, rather than requring the data of a stable, logarithmic curve. The actual moduli space of log quasimaps to more general targets is constructed in [shafi].